OpenStudy (anonymous):

Please Help!! Find an explicit rule for the nth term of the sequence. -3, -15, -75, -375, ... an = 5 • -3n + 1 an = -3 • 5n - 1 an = 5 • -3n an = -3 • 5n

OpenStudy (anonymous):

a_n = 5 • -3^n + 1 a_n = -3 • 5^n - 1 a_n = 5 • -3^n a_n = -3 • 5^n

OpenStudy (anonymous):

test one by one to see which is the correct answer let test the first one \(a_n= 5*(-3)^{n+1}\) , is it correct form of the option?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now let n =2, so a2 = 5* (-3)^3 = 5*(-27) = -135 while your second number is -15. So, the first option is not the correct one. how about the second option?? you test

OpenStudy (anonymous):

n=2 so a2=-3(5) =-15

OpenStudy (anonymous):

n=3 so a3= -3(5^2) =-75

OpenStudy (anonymous):

yup, however, to make sure of it, test for n =3

OpenStudy (anonymous):

so is the second choice the answer?

OpenStudy (anonymous):

oh, you got it. Good!! the last one, please

OpenStudy (anonymous):

should test all of the term to get 100% correct. hihihi

OpenStudy (anonymous):

a4= -3(5^3)= -375

OpenStudy (anonymous):

yay! thank you

OpenStudy (anonymous):

BiG BINGO